$SPAD/src/input WUTSET.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input generated from WuWenTsunTriangularSetXmpPage
)clear all
R := Integer
ls : List Symbol := [x,y,z,t]
V := OVAR(ls)
E := IndexedExponents V
P := NSMP(R, V)
x: P := 'x
y: P := 'y
z: P := 'z
t: P := 't
T := WUTSET(R,E,V,P)
p1 := x ** 31 - x ** 6 - x - y
p2 := x ** 8  - z
p3 := x ** 10 - t
lp := [p1, p2, p3]
characteristicSet(lp)$T
zeroSetSplit(lp)$T

———-

References^{[1] nothing}

4